# Adaptation of Agentic AI

### v2

@@ 6.2 Agent Adaptation Paradigms: A1 and A2
We now analyze the two agent-centric paradigms, which both modify the agent’s core parameters but differ fundamentally in their training signals and optimization objectives.
@@ 6.2.1 A1: Optimizing Tool Mechanics via Causal Feedback
The defining feature of A1 on-policy methods is reliance on causal, immediate, and fine-grained reward signals. The supervision source is the verifiable outcome of tool execution itself, not a downstream task metric. For example, DeepRetrieval [21] formalizes query reformulation as an MDP where reward is directly derived from retrieval metrics like Recall@K or NDCG, and RLEF [20] frames code synthesis with rewards from test-case execution. This stands in contrast to A2 signals that only evaluate the final answer.
Conceptually, A1 on-policy RL optimizes tool-use mechanics—teaching the agent how to wield tools correctly, grounding behavior in environment “physics” (“this syntax executes”, “this query retrieves”). This direct engagement with ground-truth feedback drives strong performance in domains with verifiable, deterministic outcomes.
Quantitative evidence. Mechanistic optimization under A1 achieves state-of-the-art performance in specialized domains:
@@
However, learning through trial-and-error introduces significant challenges: it requires careful reward design, KL-regularized PPO or GRPO, curriculum learning, and dynamic sampling for stable convergence.
@@ 6.2.2 A2: Optimizing Tool Strategy via Holistic Rewards
A2 methods instead use holistic, sparse, and high-level rewards based on agent output quality—typically final answer correctness—which depends on tool usage but does not directly supervise individual tool calls. ReSearch [108], trained on multi-hop QA, optimizes when and how to search. The reward asks not “was this particular search good?” but “did the entire process of thinking, searching, and reasoning lead to the correct answer?”
Thus A2 optimizes tool-use strategy or coordination. Rather than learning search mechanics (assuming a T1 retriever handles that), it learns the cognitive policy for when to search, what to search for, and how to integrate results. This strategic focus explains why ReSearch reports emergent reflection and self-correction behaviors during RL training [108].
Quantitative evidence. Strategic optimization under A2 proves highly effective for complex, multi-step reasoning:
@@
  • Factual accuracy: R1-Searcher reports up to 24% improvement over strong RAG baselines, demonstrating enhanced factual accuracy and reduced hallucination through learned retrieval policy [107].
@@
A1 & A2: Signal Source as a Reliability Axis. Beyond taxonomic categorization, the distinction between A1 and A2 fundamentally determines the granularity and scope of the adaptation signal.
  • Tool-execution signals (A1) are grounded, causal, and process-oriented. The feedback is produced by an environment or tool whose semantics are independent of the agent’s internal beliefs (e.g., code execution, retrieval metrics, formal proof checkers). This grounding enables learning that is tightly coupled to intermediate correctness and tool mastery, but often comes with higher interaction cost and environment dependence.
@@ 6.3 Tool Adaptation Paradigms: T1 and T2
We now pivot to tool-centric paradigms, which shift optimization from the expensive agent to cheaper external tools. These paradigms sacrifice some parametric flexibility (the agent policy stays fixed) but gain system-level flexibility: we can grow, specialize, and rewire the tool ecosystem without touching the main agent.
@@ 6.3.1 T1: The “Graduated Agent” as Subagent-as-Tool
T1 is defined by agent-agnostic, pre-trained, plug-and-play components. A critical concept within T1 is the subagent-as-tool, which reveals a rich development lifecycle.
@@
At the other extreme are dynamic, graduated tools: adaptive agents from §6.2 can be trained under A1 or A2 and then frozen and reused as T1 tools. This “Graduation Lifecycle” (A1 $\rightarrow$ T1) proceeds as:
@@
From the flexibility perspective, T1 offers high system-level flexibility: one can assemble different T1 tools into various configurations, or replace one tool (e.g., swap a retriever) without touching the agent. The cost of adding a capability is proportional to the size of the corresponding tool, not the backbone agent. The trade-off is that the tools are not tailored to any particular agent; the agent must adapt its prompts or orchestration logic to whatever interface the tool exposes.
6.3.2 T2: The “Symbiotic Inversion” and Subagent Federation
T2 represents a conceptual inversion. Rather than asking “how can we modify the agent to better use tools?” (A1/A2), T2 asks: “how can we modify tools to better serve a fixed agent?” This reframes the expensive foundation model from optimization target to stable supervision source.
This creates symbiotic adaptation: the frozen host agent (e.g., GPT, Claude) provides high-level reasoning and reward signals, while adaptive symbiote subagents (e.g., lightweight 7B models) learn to translate, filter, and present information in exactly the form the agent finds most useful. The core benefit is decoupling skill from knowledge. A traditional A2 agent like Search-R1 must learn (1) domain knowledge, (2) tool-use skills, and (3) task reasoning simultaneously—a complex optimization landscape. In T2, the frozen generator already possesses (1) and (3); the T2 subagent needs only learn procedural skill.
T2 subagent families also demonstrate a powerful architectural strategy: unbundling the agent’s monolithic cognitive loop (Perceive–Plan–Act–Reflect) into specialized, independently trainable submodules:
@@ 6.3.2 T2: The “Symbiotic Inversion” and Subagent Federation
  • Optimizing “Planning” (Meta-Cognitive Planners): Subagents like AI-Search Planner and AgentFlow decide how tools and specialists are deployed. AgentFlow [51] trains only a lightweight planner that orchestrates frozen specialists using trajectory-level rewards, achieving 33.1% on GAIA and surpassing $\sim$200B-parameter GPT-4.
@@
Table 5: Quantitative comparison of flagship adaptation methods.
Method Paradigm Training Signal Key Result (Quantitative) Key Insight (Qualitative)
@@
ReSearch [108] A2 Final Answer Correctness 9–22% gains over RAG Holistic RL optimizes tool strategy.
s3 [27] T2 GBR from Frozen Generator 58.9% Acc. w/ 2.4k samples Much more data-efficient than A2.
AgentFlow [51] T2 Final Answer Correctness 33.1% on GAIA (beats GPT-4) Learned orchestration of specialists.
6.4 Synthesis: The Showdown on Data Efficiency and Modularity (A2 vs. T2)
The most direct comparison arises between A2 and T2. Both aim to produce sophisticated tool-using systems, but they place the learning burden in different places. A2 adapts the agent, letting it internalize tool-use strategies; T2 adapts the tools, letting them learn to support a fixed agent.
Empirically, the contrast is stark. Comparing Search-R1 (A2) and s3 (T2), two state-of-the-art methods for retrieval-augmented generation:
  • A2 approach (Search-R1): Trains the entire agent, requiring roughly 170k examples to co-adapt internal knowledge, reasoning, and tool-use policy [108].
@@
[…]
The underlying reason is the “symbiotic inversion” discussed above. A2’s optimization landscape is high-dimensional and entangled: the agent must simultaneously adjust its knowledge, reasoning style, and tool-use policy. T2 dramatically simplifies the learning problem by assuming the backbone already solves (most of) knowledge and reasoning, and only learning a narrow procedural skill in a small subagent.
From an engineering perspective, T2 also wins on modularity. To add a new tool or update Search-R1 (A2), one must retrain the monolithic agent, potentially inducing catastrophic forgetting. In a T2 architecture, new tools can be trained and hot-swapped without touching the host agent, enabling continuous evolution of the peripheral ecosystem while the core remains stable.
@@ 6.5 Strategic Recommendations
Choosing the appropriate adaptation strategy requires balancing computational cost, data efficiency, and the need for system modularity. We organize these considerations into a strategic framework based on whether the priority lies in internal parametric adjustment or external ecosystem evolution. We detail the specific applicability, strengths, and limitations of each paradigm below:
@@
A1: Best suited for local, mechanistic mastery of verifiable tools in stable domains (e.g., retrieval, code execution, SQL). By optimizing directly on executable outcomes, A1 develops strong low-level competence and causal grounding.
  • Pros: precise control over tool behavior; robust alignment to verifiable signals.
  • Cons: high computational cost, brittle generalization, and limited transferability across tasks.
A2: Appropriate for system-level orchestration within a single agent, enabling holistic reasoning and multi-tool coordination. A2 internalizes when, how, and why to invoke tools, yielding deeply integrated reasoning patterns.
  • Pros: rich cross-tool strategies; unified end-to-end policies for complex workflows.
  • Cons: expensive monolithic retraining and susceptibility to catastrophic forgetting when scaling across domains.
@@
T1: Ideal for horizontal scalability and reusability. T1 spans both static foundational models (e.g., SAM, AlphaFold2) and “graduated” subagents—A1/A2-trained experts frozen and redeployed as reusable modules (e.g., DeepRetrieval, SWE-Grep [193]). These “subagents-as-tools” encapsulate learned procedural expertise while remaining decoupled from any specific host agent.
  • Pros: plug-and-play modularity and broad compositional flexibility across ecosystems.
  • Cons: agent-agnostic training may under-optimize tools for any particular host agent’s reasoning style.
@@
T2: Represents the symbiotic inversion: rather than adapting the agent to use tools better, T2 trains lightweight tools and subagents under frozen-agent supervision to better serve a fixed backbone (e.g., s3-style searchers, planners, advisors, and memory builders). The host agent provides high-level reasoning and reward signals, while T2 subagents learn narrow procedural skills that can be added, replaced, or composed without touching the backbone.
  • Pros: dramatic data-efficiency for new skills, mitigation of catastrophic forgetting via modular updates, and sustained compatibility with evolving or swapped host agents.
  • Cons: subagent capability is bounded by the supervising agent’s quality; multi-subagent pipelines introduce added orchestration complexity and potential error compounding.
@@
Figure 7: The $2\times 2$ adaptation landscape. The x-axis captures monolithic-to-modular evolution, while the y-axis represents local-to-systemic orchestration. A1/A2 inhabit the agent-centric half, whereas T1/T2 embody modular and system-level flexibility. Dotted arrows show that A1/A2-trained agents can graduate as tools for T1.
@@
Taken together, these four paradigms define a coherent $2\times 2$ landscape (Figure 7) along two conceptual axes: (i) the local–to–systemic spectrum (y-axis), from low-level control of specific tools (A1/T1) to holistic orchestration of multi-tool reasoning (A2/T2); and (ii) the monolithic–to–modular spectrum (x-axis), from end-to-end retraining of a single agent (A1/A2) to compositional adaptation via distributed subagents and tools (T1/T2). Viewed through this lens, A1 and A2 occupy the agent-centric half of the landscape: they directly reshape the policy parameters of the core agent, offering rich parametric flexibility but incurring heavy costs in compute, data, and stability. T1 and T2, by contrast, occupy the tool-centric half: they shift learning outward into a modular ecosystem, enabling incremental evolution, specialization, and compositional reuse. The two axes interact nonlinearly: A1 $\rightarrow$ T1 reflects the “graduation path” (frozen experts becoming reusable subagents), while A2 $\rightarrow$ T2 embodies the “federation path” (frozen backbones supervising a growing constellation of adaptive specialists). In practice, mature agentic architectures increasingly inhabit the upper-right quadrant (T2): high modularity and high orchestration, where foundation agents serve as stable cognitive centers and peripheral subagents continuously evolve to extend their capabilities.

### v3

@@ 6.2 Agent Adaptation Paradigms: A1 and A2
The two agent-centric paradigms both modify the agent’s core parameters but differ in their training signals and optimization objectives.
@@ 6.2.1 A1: Optimizing Tool Mechanics via Causal Feedback
A1 on-policy methods rely on causal, immediate, and fine-grained reward signals. The supervision source is the verifiable outcome of tool execution itself, not a downstream task metric. For example, DeepRetrieval [22] formalizes query reformulation as an MDP where reward is directly derived from retrieval metrics like Recall@K or NDCG, and RLEF [21] frames code synthesis with rewards from test-case execution. The approach contrasts with A2 signals that only evaluate the final answer.
Conceptually, A1 on-policy RL optimizes tool-use mechanics: it teaches the agent how to wield tools correctly, grounding behavior in environment “physics” (“this syntax executes”, “this query retrieves”). Direct engagement with ground-truth feedback drives strong performance in domains with verifiable, deterministic outcomes.
Quantitative evidence. Mechanistic optimization under A1 achieves strong performance in specialized domains:
@@
However, learning through trial-and-error introduces practical challenges: it requires careful reward design, KL-regularized PPO or GRPO, curriculum learning, and dynamic sampling for stable convergence.
@@ 6.2.2 A2: Optimizing Tool Strategy via Holistic Rewards
A2 methods instead use holistic, sparse, and high-level rewards based on agent output quality (typically final answer correctness) that depend on tool usage but do not directly supervise individual tool calls. ReSearch [109], trained on multi-hop QA, optimizes when and how to search. The reward asks not “was this particular search good?” but “did the entire process of thinking, searching, and reasoning lead to the correct answer?”
Thus A2 optimizes tool-use strategy and coordination. Rather than learning search mechanics (assuming a T1 retriever handles that), it learns the cognitive policy for when to search, what to search for, and how to integrate results. The strategic focus explains why ReSearch reports emergent reflection and self-correction behaviors during RL training [109].
Quantitative evidence. Strategic optimization under A2 proves effective for complex, multi-step reasoning:
@@
  • Factual accuracy: R1-Searcher reports up to 24% improvement over strong RAG baselines, with improved factual accuracy and reduced hallucination through learned retrieval policy [108].
@@
A1 & A2: Signal Source as a Reliability Axis. Beyond taxonomic categorization, the distinction between A1 and A2 determines the granularity and scope of the adaptation signal.
  • Tool-execution signals (A1) are grounded, causal, and process-oriented. The feedback is produced by an environment or tool whose semantics are independent of the agent’s internal beliefs (e.g., code execution, retrieval metrics, formal proof checkers). Such grounding enables learning that is tightly coupled to intermediate correctness and tool mastery, but often comes with higher interaction cost and environment dependence.
@@ 6.3 Tool Adaptation Paradigms: T1 and T2
Tool-centric paradigms shift optimization from the expensive agent to cheaper external tools. These paradigms sacrifice some parametric flexibility (the agent policy stays fixed) but gain system-level flexibility: the tool ecosystem can be grown, specialized, and rewired without touching the main agent.
@@ 6.3.1 T1: The “Graduated Agent” as Subagent-as-Tool
T1 is defined by agent-agnostic, pre-trained, plug-and-play components. A central concept within T1 is the subagent-as-tool, which follows a development lifecycle.
@@
At the other extreme are dynamic, graduated tools: adaptive agents from §6.2 can be trained under A1 or A2 and then frozen and reused as T1 tools. The “Graduation Lifecycle” (A1 $\rightarrow$ T1) proceeds as:
@@
From the flexibility perspective, T1 offers high system-level flexibility: different T1 tools can be assembled into various configurations, or one tool (e.g., a retriever) can be replaced without touching the agent. The cost of adding a capability is proportional to the size of the corresponding tool, not the backbone agent. The trade-off is that the tools are not tailored to any particular agent; the agent must adapt its prompts or orchestration logic to whatever interface the tool exposes.
6.3.2 T2: Inverting the Optimization Target
T2 inverts the conventional adaptation direction. Rather than adapting the agent to use tools better, T2 adapts the tools to better serve a fixed agent (see §5.2 for the full rationale). This reframes the foundation model from optimization target to supervision source.
In practice, the frozen host agent (e.g., GPT, Claude) supplies reasoning and reward signals, while lightweight subagents (e.g., 7B models) learn to reshape information for the host’s consumption. The central advantage is the decoupling of skill from knowledge. A traditional A2 agent like Search-R1 must learn (1) domain knowledge, (2) tool-use skills, and (3) task reasoning simultaneously, which creates a complex optimization landscape. In T2, the frozen generator already possesses (1) and (3); the T2 subagent needs only learn procedural skill.
T2 subagent families also instantiate an architectural strategy: unbundling the agent’s monolithic cognitive loop (Perceive-Plan-Act-Reflect) into specialized, independently trainable submodules:
@@ 6.3.2 T2: Inverting the Optimization Target
  • Optimizing “Planning” (Meta-Cognitive Planners): Subagents like AI-Search Planner and AgentFlow decide how tools and specialists are deployed. AgentFlow [52] trains only a lightweight planner that orchestrates frozen specialists using trajectory-level rewards, achieving 33.1% on GAIA and surpassing the much larger GPT-4.
@@
Table 6: Quantitative comparison of flagship adaptation methods across paradigms and domains.
Method Paradigm Domain Training Signal Key Result Key Insight
@@
RLEF [21] A1 Code Test-case pass rate Stable multi-turn PPO training Dense execution signals enable A1 RL
Code-R1 [66] A1 Code Sandboxed execution Reward quality >> data quantity Clean rewards reduce training cost
ReSearch [109] A2 RAG QA Final answer EM 9–22% gains over RAG Holistic RL optimizes tool strategy
R1-Searcher [108] A2 RAG QA Final answer correctness 24% over RAG baselines Emergent search-reason interleaving
Memento [23] T2 Memory Binary task success +4.7–9.6% on OOD tasks Memory alone can transform performance
s3 [28] T2 Retrieval GBR from frozen gen. 58.9% Acc. w/ 2.4k samples High data efficiency (see caveats in text)
AgentFlow [52] T2 Planning Final answer correctness 33.1% on GAIA (beats GPT-4) Learned orchestration of specialists
6.4 Synthesis: Data Efficiency and Modularity (A2 vs. T2)
The sharpest empirical comparison arises between A2 and T2. Both aim to produce capable tool-using systems, but they place the learning burden in different places. A2 adapts the agent, letting it internalize tool-use strategies; T2 adapts the tools, letting them learn to support a fixed agent.
The retrieval-augmented generation domain offers an illustrative case study. Comparing Search-R1 (A2) and s3 (T2):
  • A2 approach (Search-R1): Trains the entire Qwen2.5 agent, requiring roughly 170k examples to co-adapt internal knowledge, reasoning, and tool-use policy [50].
@@
Caveats. This comparison, while suggestive, is not a controlled experiment. Search-R1 trains the full Qwen2.5 agent end-to-end, whereas s3 trains only a 7B search subagent paired with a frozen generator (which may itself be Qwen2.5-7B, 14B, or Claude-3-Haiku). The two systems thus differ simultaneously in optimization target, backbone composition, and system architecture; the observed efficiency gap cannot be attributed to a single factor. Controlled cross-paradigm comparisons that isolate the effect of paradigm choice from confounding architectural differences remain an important open problem.
With this caveat, the case study illustrates a broader architectural principle. T2 simplifies the learning problem by assuming the backbone already handles most of knowledge and reasoning, and only learning a narrow procedural skill in a small subagent. A2’s optimization landscape is higher-dimensional: the agent must simultaneously adjust its knowledge, reasoning style, and tool-use policy. On specialized medical QA, T2-trained s3 reaches 76.6% accuracy vs. A2-trained Search-R1’s 71.8% [28], consistent with the hypothesis that narrower optimization targets generalize more robustly, though alternative explanations (e.g., differences in training data distribution) cannot be ruled out.
From an engineering perspective, T2 offers modularity advantages. Updating an A2 agent requires retraining the monolithic model, potentially inducing catastrophic forgetting. In a T2 architecture, new tools can be trained and hot-swapped without touching the host agent, allowing continuous evolution of the peripheral ecosystem while the core remains stable.
@@ 6.5 Strategic Recommendations
Choosing an adaptation strategy requires balancing computational cost, data efficiency, and system modularity. The following guidelines distill the empirical and architectural evidence presented above into concrete recommendations for practitioners.
@@
A1 is best suited for local, mechanistic mastery of verifiable tools in stable domains such as retrieval, code execution, and SQL. By optimizing directly on executable outcomes, A1 develops strong low-level competence and causal grounding, giving practitioners precise control over tool behavior with robust alignment to verifiable signals. The cost is high: each training run requires substantial compute, and the resulting specialization often generalizes poorly across tasks or tool interfaces.
A2 is appropriate when a single agent must orchestrate multiple tools and perform holistic reasoning. A2 internalizes when, how, and why to invoke tools, yielding deeply integrated, end-to-end policies for complex workflows. The price of this integration is expensive monolithic retraining and susceptibility to catastrophic forgetting when the agent is subsequently adapted to new domains.
@@
T1 provides horizontal scalability and reusability. The category spans both static foundational models (e.g., SAM, AlphaFold2) and “graduated” subagents, A1/A2-trained experts that are frozen and redeployed as reusable modules (e.g., DeepRetrieval, SWE-Grep [243]). These “subagents-as-tools” encapsulate learned procedural expertise while remaining decoupled from any specific host agent, enabling plug-and-play modularity and broad compositional flexibility. Because T1 tools are trained without reference to a particular agent, they may be under-optimized for any given host’s reasoning style.
@@
T2 inverts the adaptation question: rather than adapting the agent to use tools better, it trains lightweight tools and subagents under frozen-agent supervision to better serve a fixed backbone (e.g., s3-style searchers, planners, advisors, and memory builders). The host agent provides high-level reasoning and reward signals, while T2 subagents learn narrow procedural skills that can be added, replaced, or composed without touching the backbone. T2 achieves high data efficiency for new skills and mitigates catastrophic forgetting through modular updates, but subagent capability is bounded by the supervising agent’s quality, and multi-subagent pipelines introduce orchestration complexity and potential error compounding.
@@
Figure 7: An interpretive projection of the four adaptation paradigms. Unlike the definitional axes in §3 (what is adapted $\times$ signal source), this figure offers a complementary design-space view. The x-axis captures monolithic-to-modular evolution, while the y-axis represents local-to-systemic orchestration. A1/A2 inhabit the agent-centric half, whereas T1/T2 represent modular and system-level flexibility. Dotted arrows show that A1/A2-trained agents can graduate as tools for T1.
@@
Taken together, these four paradigms can be projected onto a complementary $2\times 2$ design-space view (Figure 7). Note that this projection is an interpretive comparison, not the definitional axes introduced in §3; the definitional taxonomy classifies methods by what is adapted and how the signal is obtained, whereas the axes here capture broader architectural tendencies: (i) the local-to-systemic spectrum (y-axis), from low-level control of specific tools (A1/T1) to holistic orchestration of multi-tool reasoning (A2/T2); and (ii) the monolithic-to-modular spectrum (x-axis), from end-to-end retraining of a single agent (A1/A2) to compositional adaptation via distributed subagents and tools (T1/T2). Viewed through this lens, A1 and A2 occupy the agent-centric half of the landscape: they directly reshape the policy parameters of the core agent, offering rich parametric flexibility but incurring heavy costs in compute, data, and stability. T1 and T2, by contrast, occupy the tool-centric half: they shift learning outward into a modular ecosystem, permitting incremental evolution, specialization, and compositional reuse. The two axes interact nonlinearly: A1 $\rightarrow$ T1 reflects the “graduation path” (frozen experts becoming reusable subagents), while A2 $\rightarrow$ T2 follows the “federation path” (frozen backbones supervising a growing constellation of adaptive specialists). In practice, mature agentic architectures increasingly inhabit the upper-right quadrant (T2): high modularity and high orchestration, where foundation agents serve as stable cognitive centers and peripheral subagents continuously evolve to extend their capabilities.